OpenStudy (anonymous):

A running man has half the kinetic energy of a boy half his mass. The man speeds up by 1.70 m/s and then has the same kinetic energy as the boy. What is the speed (in meters/second) of the boy? I feel like this question is lacking information, but clearly its not since its a homework question. Any hints would be really appreciated.

OpenStudy (irishboy123):

you need to build a few equations. (1) if the KE of the man to start with = 1/2 M V^2 and the boy runs at U, we know what? we can write an equation in V, U and M. recap, the boy's mass is M/2.

OpenStudy (anonymous):

Ohhh, I see. I did that, but I wasn't multiplying my answer of 4.106 by 2 to get the answer for the boys mass, which is 8.21. Thank you @IrishBoy123 !

OpenStudy (irishboy123):

if you want to check answer, tell me and i'll work it out too...

OpenStudy (anonymous):

Sure, I was doing: Man's velocity = v Boys velocity = V Mass of the man = M \[\frac{ 1 }{ 2 } mv ^{2} = (\frac{ 1 }{ 2 } )(\frac{ 1 }{ 2 }) mV ^{2}\] \[v ^{2} = 0.25V ^{2}\] \[\frac{ 1 }{ 2 } M (v+1.7) ^{2} = (\frac{ 1 }{ 2 }) (\frac{ 1 }{ 2 }) MV ^{2}\] (\[(\frac{ (v+1.7) }{ v })^{2} = 2V\] \[v + 1.7 = 1.414v\] \[0.414v = 1.7\] \[v=\frac{ 1.7 }{ 0.414 } = 4.106\] \[4.106 \times 2 = 8.21 (answer)\] I know the answer is right because our homework is a computer system that immediately tells us. :D Sorry if I skipped a lot of algebra.